OpenStudy (aonz):

Help me pls!!! my coaching teacher is going 2 kill me

OpenStudy (aonz):

\[\frac{ 2 }{ 9 } \] of the people on a restaurant are adults. if there are 95 more children than adults, how many children are there in the restaurant

OpenStudy (kuoministers):

he really gonna kill u???

OpenStudy (aonz):

yes

OpenStudy (kuoministers):

i would be dead scared ^^

OpenStudy (aonz):

pls help

OpenStudy (anonymous):

Okay I'll try :)

OpenStudy (kuoministers):

7/9 people are children

OpenStudy (aonz):

i know but in number?

OpenStudy (kuoministers):

how would i know.. @FirstFrostByte is pro

OpenStudy (aonz):

uhhhhh frost bite?

OpenStudy (anonymous):

95 more children, so 7/9 of the people the children portion is 95. Knowing that: (7/9 - 2/9)*x = 95 5/9 * x = 95 So the total people would be: 95 * 9/5 = 171 Next we do: 2/9 *171 = 38 To get the adult count. And this for the children count: 7/9 *171 = 133

OpenStudy (kuoministers):

im pretty sure u lost him

OpenStudy (anonymous):

There are multiple ways to solve this... I don't know which method he wants...

OpenStudy (aonz):

well i kinda like that method u just did

OpenStudy (anonymous):

@AonZ You know what a variable is right?

OpenStudy (aonz):

not really

OpenStudy (anonymous):

Okay let me do this step by step.

OpenStudy (aonz):

ty

OpenStudy (anonymous):

\[9/9\]would be everyone in the restaurant. Given to us is 2/9 is adults. So that leaves? 9 - 2 = 7 7/9 Must be children. Do you understand so far before I continue?

OpenStudy (aonz):

yes

OpenStudy (aonz):

100%

OpenStudy (anonymous):

For each "9" people there are 5 more then adults. So: Let's use a variable now which will be "P" for people. 5/9 * P = 95 Now let's flip over 95 to the other side. 95 * 9/5 (it becomes 9/5 due to a rule) = P 95 * 9/5 = 171 Correct? Now we know the total amount of people. Do you understand?

OpenStudy (aonz):

ohhhhh i get now

OpenStudy (aonz):

i get the question

OpenStudy (aonz):

it is just that part that i got lost

OpenStudy (aonz):

yay so their is 133 children

OpenStudy (anonymous):

Okay. So: Since we now know the total: Next we do: 2/9(Adults) *171 = 38 Adults And this for the children count: 7/9(Children) *171 = 133 Children So there for 133 Children are in the restaurant.

OpenStudy (aonz):

and 38 adults

OpenStudy (aonz):

ty frost byte my coaching techer will not kill me now

OpenStudy (aonz):

done my homework

OpenStudy (anonymous):

Haha :) I'm happy you didn't die. There is another method you could do: Take 95 / 5(Since we know there is 5 more children then adults.) And you get 19. So you need 19 sets of 2 adults and 7 children to equal the total. 19 x 2 = 38 Adults 19 x 7 = 133 Children Check your work: 133 - 38 = 95 See?

OpenStudy (aonz):

ohhhh i think this is a better method for me

OpenStudy (anonymous):

I guess. But if your in Pre-Algebra. Go with the first method. I'm cleaner and faster. And you'll use it more often too.

OpenStudy (aonz):

kk cya